OpenStudy (anonymous):

find P if I=1260, R=4.5, and T=7. I got 1260=P*4.5*7 and than I got 1260=P*31.5. So what would I do next

OpenStudy (ranga):

Simple Interest I = P*R*T R is the rate of interest and it should be in decimal not percentage. R = 4.5% = 0.045 1260 = P * .045 * 7 1260 = P * 0.315 Divide both sides by 0.315 1260/0.315 = P P = ?